# Max Norris1, Kobi Gal1 and Sahan Bulathwela2

### v1

@@ Introduction
Modelling latent variables such as student knowledge (Bulathwela et al. 2020; Qiu et al. 2024), skills and learning progression is a core challenge in the design of Student Models (Mousavinasab et al. 2021), that make up a key part of intelligent learning companions of the future (Perez-Ortiz et al. 2021). Knowledge Tracing (KT) addresses this by using students’ responses to questions—typically, whether they answered correctly—as observable signals of their underlying knowledge state. A wide range of KT models have been developed to predict student performance on future questions (Khajah et al. 2014; Cen, Koedinger, and Junker 2006; Pavlik, Cen, and Koedinger 2009; Pandey and Karypis 2019). More recent approaches leverage deep learning to model complex patterns in student interaction histories, enabling more accurate response prediction (Piech et al. 2015; Liu et al. 2019).
Deep learning approaches to Knowledge Tracing (KT) have demonstrated considerable potential, yet their progress is limited by dataset constraints. Many widely used open-source KT datasets (e.g., ASSISTments, EdNet (Feng, Heffernan, and Koedinger 2009; Choi et al. 2020)) exclude the text of the questions answered by students (Choi et al. 2020), a common omission in large datasets released by private companies due to intellectual property restrictions.
@@
The text of a question often carries rich pedagogical signals that could enhance predictive accuracy in Knowledge Tracing models. For example, Barbu et al. (2010) found that linguistic complexity affects student comprehension in math problem-solving, showing how textual features influence performance in real world settings (Barbu and Beal 2010). More broadly, full question text carries finer-grained information about topics, structure, and wording that simple categorical labels fail to capture. Reducing the question to these coarser-grain annotations discards this nuance, limiting the model’s ability to access useful signals or even learn new ones. Natural language is also helpful when learning generalizable representations in the question space; this has benefits for cold-start scenarios where students encounter unseen questions—a challenge for current KT systems (Jung et al. 2024).
We hypothesise that effectively incorporating question text alongside interaction history metadata can lead to improved knowledge tracing performance and better cold-start generalisation. To this end, we propose a novel approach that tightly integrates question text into the KT objective—what we term Next Token Knowledge Tracing (NTKT). Leveraging the strengths of LLMs in natural language understanding (Jin et al. 2023; Singhal et al. 2023; Zhao et al. 2023), NTKT reformulates knowledge tracing as a next-token prediction task. Student interaction histories are converted into natural language sequences, allowing fine-tuned LLMs to predict the next token—i.e., the student’s response—based on their full contextual history, including question text.
To validate our approach, we conduct three experiments. First, we benchmark NTKT against existing Deep Learning baselines, demonstrating improved predictive accuracy. Second, we conduct controlled ablation studies to isolate the impact of textual information by comparing model performance across varying input representations—from full question text to ID-only formats. Finally, we assess NTKT’s generalisation capabilities under two cold-start scenarios, 1) user cold start and 2) question cold start.
@@
Early KT approaches used statistical models such as Bayesian Knowledge Tracing (BKT) (Khajah et al. 2014; Pardos and Heffernan 2010; Yudelson, Koedinger, and Gordon 2013), which represent knowledge states as binary variables and offer interpretability and simplicity. Other models like Learning Factor Analysis (LFA) and Performance Factor Analysis (PFA) (Cen, Koedinger, and Junker 2006; Pavlik, Cen, and Koedinger 2009) extend this approach using logistic regression to model student performance and learning.
Deep learning has since reshaped KT. Deep Knowledge Tracing (DKT) (Piech et al. 2015) introduced RNNs for sequence modelling. Following this work was a range of extensions starting with models such as DKVMN (Zhang et al. 2017), which incorporated memory networks. Next, a range of attention-based models, including SAKT (Pandey and Karypis 2019), AKT (Ghosh, Heffernan, and Lan 2020) and RKT (Pandey and Srivastava 2020), demonstrated improved performance on long sequences. Other recent state-of-the-art methods include the DTransformer (Yin et al. 2023), which uses a contrastive loss for stable knowledge tracing.
@@
More recently, LLMs have been applied to KT via prompting (Neshaei et al. 2024), multi-agent setups (Trifa, Hedhili, and Chaari 2019), and fine-tuning on partial student histories (Jung et al. 2024). These approaches undermine the value of domain-adaptation of content representations, whereas our work shows the performance gain through finetuning of LLMs to KT.
The language model–based knowledge tracing method (LKT) has been shown to apply smaller encoder-based pre-trained models effectively to KT in a masked language modelling setup (Lee et al. 2024). While this method boasts impressive performance, it is limited by short context windows (512 tokens) —capturing only a few student interactions at a time—despite students often having long interaction histories. Furthermore, their experiments are based on datasets not widely used in the KT community. Crucially, this method relies on smaller encoder-based models with a masked language modelling objective, in contrast to the more recent advancements with much larger decoder-only architectures.
@@
We focus on the Knowledge Tracing (KT) scenario where a learner $\ell$ in the learner population of L{L} learners, interacts with a series of exercises sequentially in discrete timesteps $1\dots t\dots T_{\ell}$. At timestep $t$, learner $\ell$ attempts the multiple-choice question exercise $e^{\ell}_{t}$ that consists of a question $q^{\ell}_{t}$ and answer options $O^{\ell}_{t}\in\{o^{\ell}_{t_{1}}\dots o^{\ell}_{t_{j}}\dots o^{\ell}_{t_{J}}\}$ leading to the answer outcome $a^{\ell}_{t}\in\{1,0\}$. The modelling task entails predicting the answer outcome $\hat{a}^{\ell}_{t}$ using the learner’s exercise answering history $\mathcal{H}^{\ell}_{t}\in\{e^{\ell}_{1}\dots e^{\ell}_{h}\dots e^{\ell}_{t-1}\}$ where each exercise $e^{\ell}_{h}$ has a question $q^{\ell}_{h}$ and answer options $O^{\ell}_{h}$. Formally, the modelling problem can be outlined as per expression 1.
$P(a^{\ell}_{t}=1|q^{\ell}_{t},O^{\ell}_{t})P(\mathcal{H}^{\ell}_{t})\text{ where }P(\mathcal{H}^{\ell}_{t})=\prod_{h=1}^{t-1}P(q^{\ell}_{h},O^{\ell}_{h},a^{\ell}_{h})$ (1)
@@
We evaluated our approach against several established baselines used for knowledge tracing. We modified the AKT baseline model to incorporate question content embeddings and investigated whether retrofitting existing architectures with textual information could improve their performance.
DKT
is a deep learning knowledge tracing model based on (Piech et al. 2015) that utilises an LSTM to capture temporal dependencies in student learning trajectories and predict future performance. The model maintains a hidden state $h^{\ell}_{h}$ representing the student’s knowledge at time $h$. At each step, this hidden state is updated using an LSTM with the current interaction features:
$h^{\ell}_{h+1}=\text{LSTM}(h^{\ell}_{h},x^{\ell}_{h})$ (2)
where $x^{\ell}_{h}$ encodes the interaction features, including the question ID embedding $\text{Eq}(q^{\ell}_{h})$ and previous response. The model predicts response probabilities by applying a sigmoid function to a linear projection of the hidden state:
$P(a^{\ell}_{t}=1\mid h^{\ell}_{t})=\sigma(Wh^{\ell}_{t}+b)$ (3)
AKT
(Attentive Knowledge Tracing) is a self-attention based model (Ghosh, Heffernan, and Lan 2020) that uses multi-head attention mechanisms to capture relationships between interactions. The model employs transformer-style attention blocks with exponential decay to weight recent interactions more heavily.
@@
extends the standard AKT by incorporating semantic embeddings of question content. Question text is embedded using pre-trained sentence transformers (all-MiniLM-L6-v2) to generate 384-dimensional embeddings that capture semantic meaning beyond question IDs. These embeddings are integrated into the model through a projection layer:
$x^{\ell,\text{aug}}_{t}=[x^{\ell}_{t}\oplus\text{proj}(\text{embed}(q^{\ell,\text{text}}_{t}))]$ (4)
where $\text{embed}(q_{t}^{\ell,\text{text}})$ represents the pre-computed embedding of the question text, and $\text{proj}(\cdot)$ is a learned projection to align the embedding dimension with the model’s hidden dimension. The attention mechanisms then operate on these augmented representations.
@@
is a transformer-based architecture specifically designed for stable knowledge tracing. The Dtransformer uses an attention mechanism in conjuction with contrastive learning objectives to obtain stable and consistent knowledge states. The model achieves state of the art scores across major KT benchmarks (Yin et al. 2023).
@@
<history>:
<Q>: [Question 1] <options>: [Options A-D] </options> <QID>[ID]</QID> <C>[Concept]</C> </Q><cr>[Correct/Incorrect]</cr>
<Q>: [Question N] <options>: [Options A-D] </options> [...additoinal metadata]</Q><cr>[Correct/Incorrect]</cr>
</history>
What do you predict they will answer for the target question:
<target>[Target Question] <options>
[Option A-D]</options> </target>:
@@
To assess the effect of model scale, we fine-tuned several open-source decoder-only LLMs from the LLaMA 3.1/3.2 families, specifically 1B, 3B, and 8B parameter variants. These models serve as the foundation for NTKT, fine-tuned using the selective masking objective described earlier. The best-performing configuration from this stage is used in the subsequent ablation and cold-start analyses to ensure computational efficiency.
@@
Among the evaluated configurations, the LLaMA 3B NTKT model achieves the best overall performance (F1 = 90.20%, Accuracy = 86.70%, AUC = 95.72%), followed closely by the 8B and 1B variants. These results demonstrate that pretrained language model representations can be effectively adapted to model student learning trajectories, even at relatively small model scales.
Compared with strong neural baselines such as DKT, AKT, and DTransformer, NTKT shows substantial gains—exceeding 10 percentage points in both F1 and AUC. This indicates that incorporating question text and leveraging pretrained representations provides richer contextual understanding of student–question interactions, enabling more accurate prediction of student performance.
@@
Table 1: Performance (mean ± std) of different model families on the Binary Knowledge Tracing task (Eedi dataset). Metrics include F1, Accuracy, and AUC. The best and next best performances are shown in bold and italic, respectively.
Model F1 (%) Accuracy (%) AUC (%)
Baseline Models
No-FT 70.52 ± .0000 57.20 ± .0000 61.20 ± .0000
DKT 77.46 ± .0000 63.22 ± .0000 73.23 ± .0013
AKT 76.92 ± .0043 68.31 ± .0010 72.37 ± .0013
DTransformer 76.80 ± .0020 68.88 ± .0012 73.23 ± .0076
AKTtext 77.02 ± .0008 68.90 ± .0004 72.23 ± .0031
@@
Table 2 shows how varying levels of textual information influence NTKT performance across different model sizes. In all cases, including the full question text yields the best results, confirming that natural language content provides essential pedagogical cues that enhance predictive accuracy and calibration.
The concept-only configuration, which retains semantic tags but omits full text, leads to a moderate decline across all metrics. Meanwhile, the ID-only condition—where models receive only symbolic question identifiers—retains competitive F1 and Accuracy but exhibits noticeable AUC degradation, suggesting less reliable probability estimates.
@@
Table 2: Impact of input representation on NTKT performance across different text information levels and model sizes. The best and next best performances for each model are shown in bold and italic, respectively.
Model & Input F1 (%) Acc (%) AUC (%)
NTKT (LLaMA3.2-1B)
   ID-only 87.37 ± .0114 84.57 ± .0120 85.39 ± .0296
   Concept-only 88.13 ± .0103 84.88 ± .0131 85.63 ± .0256
   Full Text 88.28 ± .0143 85.11 ± .0244 93.35 ± .0183
NTKT (LLaMA3.2-3B)
   ID-only 86.55 ± .0131 83.72 ± .0131 85.39 ± .0296
   Concept-only 88.34 ± .0103 85.52 ± .0131 85.63 ± .0256
   Full Text 90.20 ± .0013 86.70 ± .0020 95.72 ± .0018
NTKT (LLaMA3.1-8B)
   ID-only 85.97 ± .0553 80.67 ± .1078 86.81 ± .0184
   Concept-only 87.25 ± .0221 83.34 ± .0389 90.60 ± .0056
   Full Text 88.37 ± .0043 85.74 ± .0040 93.53 ± .0091
@@
Cold-start robustness for unseen questions is shown in Figure 3. Baseline models exhibit a clear performance drop between seen and unseen questions (0.777 to 0.732, $p<0.001$), whereas NTKT maintains stable performance (0.843 on both, $p>0.5$). Two of the three NTKT models even show slight gains on unseen items. This demonstrates strong generalisation to new content, driven by NTKT’s ability to leverage pretrained representations and linguistic cues rather than memorised identifiers.
@@
Several promising directions exist for future research. These include investigating cross-domain adaptation to other subject areas and multi-lingual settings. The framework could also be expanded to handle multi-modal content like mathematical equations and diagrams for STEM education. Another key direction is developing interpretability tools to extract actionable pedagogical insights from attention patterns as well as investigating sources of potential bias within NTKT so that the outputs can be presented to learners via intelligent user interfaces for learning (e.g. like X5Learn (Pérez Ortiz et al. 2022)). Furthermore, NTKT’s generative capabilities could be leveraged to provide personalized questions (Li, Cukurova, and Bulathwela 2025) and explanations based on predicted behavior. Future work should also explore computational optimizations to make NTKT more accessible for resource-constrained educational environments while maintaining its performance advantages.

### v2

@@ Introduction
Modelling latent variables such as student knowledge (Bulathwela et al. 2020; Qiu et al. 2024), skills and learning progression is a core challenge in the design of Student Models (Mousavinasab et al. 2021), that make up a key part of intelligent learning companions of the future (Perez-Ortiz et al. 2021). Knowledge Tracing (KT) addresses this by using students’ responses to questions, as observable signals of their underlying knowledge state. A wide range of KT models have been developed to predict student performance on future questions (Khajah et al. 2014; Cen et al. 2006; Pavlik et al. 2009; Pandey and Karypis 2019). More recent approaches leverage deep learning to model complex patterns in student interaction histories, enabling more accurate response prediction (Piech et al. 2015; Liu et al. 2019).
Deep learning approaches to Knowledge Tracing (KT) have demonstrated considerable potential, yet their progress is limited by dataset constraints. Many widely used open-source KT datasets (e.g., ASSISTments, EdNet (Feng et al. 2009; Choi et al. 2020)) exclude the text of the questions answered by students (Choi et al. 2020), a common omission in large datasets released by private companies due to intellectual property restrictions.
@@
The text of a question often carries rich pedagogical signals that could enhance predictive accuracy in Knowledge Tracing models. For example, Barbu et al. (2010) found that linguistic complexity affects student comprehension in math problem-solving, showing how textual features influence performance in real world settings (Barbu and Beal 2010). More broadly, full question text carries finer-grained information about topics, structure, and wording that simple categorical labels fail to capture. Reducing the question to these coarser-grain annotations discards this nuance, limiting the model’s ability to access useful signals or even learn new ones. Natural language is also helpful when learning generalizable representations in the question space; this has benefits for cold-start scenarios where students encounter unseen questions, a challenge for current KT systems (Jung et al. 2024).
We hypothesise that effectively incorporating question text alongside interaction history metadata can lead to improved knowledge tracing performance and better cold-start generalisation. To this end, we propose a novel approach that tightly integrates question text into the KT objective: what we term Next Token Knowledge Tracing (NTKT). Leveraging the strengths of LLMs in natural language understanding (Jin et al. 2023; Singhal et al. 2023; Zhao et al. 2023), NTKT reformulates knowledge tracing as a next-token prediction task. Student interaction histories are converted into natural language sequences, allowing fine-tuned LLMs to predict the next token i.e., the student’s response—based on their full contextual history, including question text.
To validate our approach, we conduct three experiments. First, we benchmark NTKT against existing Deep Learning baselines, demonstrating improved predictive accuracy. Second, we conduct controlled ablation studies to isolate the impact of textual information by comparing model performance across varying input representations, from full question text to ID-only formats. Finally, we assess NTKT’s generalisation capabilities under two cold-start scenarios, 1) user cold start and 2) question cold start.
@@
Early KT approaches used statistical models such as Bayesian Knowledge Tracing (BKT) (Khajah et al. 2014; Pardos and Heffernan 2010; Yudelson et al. 2013), which represent knowledge states as binary variables and offer interpretability and simplicity. Other models like Learning Factor Analysis (LFA) and Performance Factor Analysis (PFA) (Cen et al. 2006; Pavlik et al. 2009) extend this approach using logistic regression to model student performance and learning.
Deep learning has since reshaped KT. Deep Knowledge Tracing (DKT) (Piech et al. 2015) introduced RNNs for sequence modelling. Following this work was a range of extensions starting with models such as DKVMN (Zhang et al. 2017), which incorporated memory networks. Next, a range of attention-based models, including SAKT (Pandey and Karypis 2019), AKT (Ghosh et al. 2020) and RKT (Pandey and Srivastava 2020), demonstrated improved performance on long sequences. Other recent state-of-the-art methods include the DTransformer (Yin et al. 2023), which uses a contrastive loss for stable knowledge tracing.
@@
More recently, LLMs have been applied to KT via prompting (Neshaei et al. 2024), multi-agent setups (Trifa et al. 2019), and fine-tuning on partial student histories (Jung et al. 2024). These approaches undermine the value of domain-adaptation of content representations, whereas our work shows the performance gain through finetuning of LLMs to KT.
The language model–based knowledge tracing method (LKT) has been shown to apply smaller encoder-based pre-trained models effectively to KT in a masked language modelling setup (Lee et al. 2024). While this method boasts impressive performance, it is limited by short context windows (512 tokens), capturing only a few student interactions at a time—despite students often having long interaction histories. Furthermore, their experiments are based on datasets not widely used in the KT community. Crucially, this method relies on smaller encoder-based models with a masked language modelling objective, in contrast to the more recent advancements with much larger decoder-only architectures.
@@
Each learner $\ell\in\{1,\dots,L\}$ interacts with a sequence of exercises over discrete timesteps $t=1,\dots,T_{\ell}$. At timestep $t$, learner $\ell$ attempts a multiple-choice exercise $e_{t}^{\ell}=(q_{t}^{\ell},O_{t}^{\ell})$, where $q_{t}^{\ell}$ denotes the question and $O_{t}^{\ell}=\{o_{t_{1}}^{\ell},\dots,o_{t_{J}}^{\ell}\}$ is the set of answer options. The learner’s binary correctness outcome is denoted $a_{t}^{\ell}\in\{0,1\}$. We denote the learner’s interaction history up to time $t$ as
$\mathcal{H}_{t}^{\ell}\;=\;\bigl\{(e_{1}^{\ell},a_{1}^{\ell}),\dots,(e_{t-1}^{\ell},a_{t-1}^{\ell})\bigr\}.$ (1)
Following the standard knowledge tracing formulation, the objective at timestep $t$ is to predict whether learner $\ell$ will answer the current exercise correctly. Given the current exercise $e_{t}^{\ell}=(q_{t}^{\ell},O_{t}^{\ell})$ and the historical sequence $\mathcal{H}_{t}^{\ell}$, the predictive model estimates
$P_{\Theta}\!\left(a_{t}^{\ell}=1\;\middle|\;q_{t}^{\ell},O_{t}^{\ell},\mathcal{H}_{t}^{\ell}\right).$ (2)
where $P_{\Theta}$ denotes the model parameterized by $\Theta$.
@@
To assess the performance of NTKT, we compare it against several established knowledge tracing architectures. These baselines represent the evolution of the field from recurrent neural networks to modern attention-based systems. We specifically adapted certain models to incorporate textual data to ensure a fair comparison with our text-aware approach.
Deep Knowledge Tracing (DKT)
(Piech et al. 2015) DKT utilizes Long Short-Term Memory (LSTM) networks to process a student’s interaction history as a chronological sequence. The model maintains a compressed numerical representation, or hidden state, of a student’s latent knowledge, which is updated at each timestep based on the specific question identifier and the student’s previous correctness. The final output is a probability indicating the likelihood of the student answering the next question correctly.
Attentive Knowledge Tracing (AKT)
(Ghosh et al. 2020) shifts away from recurrent architectures in favor of a self-attention mechanism. This model uses a Transformer-style approach to calculate the relevance between a student’s past interactions and the current target question. Unlike recurrent models that compress history into a single cumulative state, AKT can directly ”attend” to specific past experiences. It incorporates a temporal decay function to prioritize recent learning while using specialized embeddings to capture the relationship between different educational concepts.
@@
is our augmented version of the AKT model, designed to evaluate if traditional KT architectures can be effectively retrofitted with access to full question text. We use a pre-trained sentence transformer (all-MiniLM-L6-v2) to convert the question text into a dense 384-dimensional numerical vector. This embedding is then projected and concatenated with the standard question and concept identifiers. This allows the AKT attention mechanism to factor in the semantic meaning of the question text alongside categorical metadata.
@@
(Yin et al. 2023) is a state-of-the-art Transformer-based architecture that introduces more stable knowledge tracking through a decoupled representation of questions and student proficiency. It employs a contrastive learning objective to ensure that similar learning states are mapped closely together in the model’s internal representation space.
No-FT (Zero-Shot LLaMA)
serves as a control without domain-specific adaptation. We use the LLaMA 3.2 3B model in its base instructional state, providing it with the same history and target question as our NTKT approach but without any fine-tuning on student interaction data. This allows us to isolate the performance gains attributable specifically to our fine-tuning framework.
@@
To assess the effect of model scale, we fine-tuned several open-source decoder-only LLMs from the LLaMA 3.2 family, specifically 1B and 3B parameter variants. These models serve as the foundation for NTKT, fine-tuned using the selective masking objective described earlier. The best-performing configuration from this stage is used in the subsequent ablation and cold-start analyses to ensure computational efficiency.
@@
Among the evaluated configurations, the LLaMA 3B NTKT model achieves the best overall performance (F1 = 85.22%, Accuracy = 78.87%, AUC = 90.32%), followed by the 1B variant. These results demonstrate that pretrained language model representations can be effectively adapted to model student learning trajectories, even at relatively small model scales.
Compared with strong neural baselines such as DKT, AKT, and DTransformer, NTKT shows substantial gains—exceeding 10 percentage points in AUC over the best-performing baseline. This indicates that incorporating question text and leveraging pretrained representations provides richer contextual understanding of student–question interactions, enabling more accurate prediction of student performance.
@@
Table 1: Predictive Performance on the Eedi Dataset. Results show Mean $\pm$ SD across 5 runs. Best results are bolded.
Model F1 Accuracy AUC
Baselines
DKT 77.46 $\pm$ .0000 63.22 $\pm$ .0000 73.23 $\pm$ .0013
AKT 76.92 $\pm$ .0043 68.31 $\pm$ .0010 72.37 $\pm$ .0013
AKTtext 77.02 $\pm$ .0008 68.90 $\pm$ .0008 72.23 $\pm$ .0031
DTransformer 76.80 $\pm$ .0020 68.88 $\pm$ .0012 73.23 $\pm$ .0076
No-FT 74.25 $\pm$ .0030 66.78 $\pm$ .0020 75.60 $\pm$ .0030
@@
Table 2 shows how varying levels of textual information influence NTKT performance. In all cases, including the full question text yields the best results, confirming that natural language content provides essential pedagogical cues that enhance predictive accuracy and calibration.
The concept-only configuration, which retains semantic tags but omits full text, leads to a decline across all metrics (e.g., AUC dropping from 90.32% to 84.59%). Meanwhile, the ID-only condition, where models receive only symbolic question identifiers, retains competitive F1 but exhibits significant AUC degradation (77.06%), suggesting less reliable probability estimates when textual context is removed.
@@
Table 2: Eedi NTKT (LLaMA 3B) performance using different feature representations. Best and second-best results are shown in bold and italic, respectively.
Features F1 Accuracy AUC
ID-only 79.16 ± .0131 69.22 ± .0120 77.06 ± .0085
Concept-only 83.22 ± .0114 76.77 ± .0105 84.59 ± .0062
Full Text 85.22 ± .0013 78.87 ± .0020 90.32 ± .0014
@@
Cold-start robustness for unseen questions is shown in Figure 3. Baseline models exhibit a clear performance drop between seen and unseen questions (0.777 to 0.732, $p<0.001$), whereas NTKT maintains stable performance (0.843 on both, $p>0.5$). Both NTKT model variants show stable or slightly improved gains on unseen items. This demonstrates strong generalisation to new content, driven by NTKT’s ability to leverage pretrained representations and linguistic cues rather than memorised identifiers.
@@
Several promising directions exist for future research. These include investigating cross-domain adaptation to other subject areas and multi-lingual settings. The framework could also be expanded to handle multi-modal content like mathematical equations and diagrams for STEM education. Another key direction is developing interpretability tools to extract actionable pedagogical insights from attention patterns as well as investigating sources of potential bias within NTKT so that the outputs can be presented to learners via intelligent user interfaces for learning (e.g. like X5Learn (Pérez Ortiz et al. 2022)). Furthermore, NTKT’s generative capabilities could be leveraged to provide personalized questions (Li et al. 2025) and explanations based on predicted behavior. Future work should also explore computational optimizations to make NTKT more accessible for resource-constrained educational environments while maintaining its performance advantages.
Appendix A Prompt Templates and Implementation Details
This appendix provides the full textual structure used to prompt the Large Language Model for the Next Token Knowledge Tracing (NTKT) task.
<history>:
<Q>
<text>[Question 1]</text>
<options>[Options A-D]</options>
<QID>[ID]</QID>
<C>[Concept]</C>
</Q><cr>[Correct/Incorrect]</cr>
<Q>
<text>[Question N]</text>
<options>[Options A-D]</options>
<QID>[ID]</QID>
<C>[Concept]</C>
</Q><cr>[Correct/Incorrect]</cr>
</history>
What do you predict they will answer for the target question:
<target>
<text>[Target Question]</text>
<options>[Option A-D]</options>
<QID>[ID]</QID>
<C>[Concept]</C>
</target>: